OpenStudy (anonymous):

Use the distributive property to simplify the expression below: −3(3d − 1) + 5d

OpenStudy (anonymous):

−4d + 3 −4d − 3 −d + 3 −d − 3

OpenStudy (anonymous):

well, first off what is the distributive property?

OpenStudy (anonymous):

the distributive property says that if you have numbers like this: \[a(b+c)\] then you can rewrite them like this:\[ab+ac\]

OpenStudy (anonymous):

basically you take what is on the outside of the parenthesis and multiply it by each thing inside the parenthesis. like this:

OpenStudy (anonymous):

so what we get when we use the distributive property on this equation is this: \[-3(3d - 1) + 5 = -9d+3+5\]

OpenStudy (anonymous):

-3(3d-1)+5d=-9d-3-15d

OpenStudy (anonymous):

i mean -24d+3

OpenStudy (anonymous):

combined like terms

OpenStudy (anonymous):

sorry, forgot a "d" so when you finally get down to it and combine the like terms what you are going to get is: \[-9d+5d+3 = -4d+3\]

OpenStudy (anonymous):

so the correct answer is \[-4d+3\]

OpenStudy (anonymous):

the small mistake that @pudgesan made was that when he multiplied his -3, he didn't notice that the +5d was outside of the parenthesis and therefore shouldnt have gotten multiplied by the -3

OpenStudy (anonymous):

i am a girl

OpenStudy (anonymous):

oops! sorry!!

OpenStudy (anonymous):

the small mistake that @pudgesan made was that when SHE multiplied his -3, SHE didn't notice that the +5d was outside of the parenthesis and therefore shouldnt have gotten multiplied by the -3

OpenStudy (anonymous):

it is okay

OpenStudy (anonymous):

better? ^_^

OpenStudy (anonymous):

does that make sense @ahoward79 ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Yes! Thank you so much! You are an AMAZING teacher!